OpenStudy (rissyroo118):

help one these 2 questions please!! Jose received $30 for his birthday and put it in his piggy bank. Each week, he puts 4 more dollars in his bank. The amount of money in dollars D in his bank is given by D = 4w + 30 where w is the number of weeks he has saved 4 dollars. His sister Marisol makes a table of reasonable values for w. Marisol selects 16 as a values for w. What is the corresponding table entry for D? (Points : 1) 34 434 50 94

OpenStudy (rissyroo118):

Andre received $60 for his birthday and put it in his piggy bank. Each week, he puts 3 more dollars in his bank. The amount of money in dollars D in his bank is given by D = 3w + 60 where w is the number of weeks he has saved 3 dollars. His sister Lana makes a table of reasonable values for w. Lana selects 42 as a value for w. What is the corresponding table entry for D? (Points : 1) 54 105 186 402

OpenStudy (anonymous):

1st Question: D = 4w + 30 = 4(16) + 30 =94 2nd Question: 186 same method